# Deep Multimodal Learning with Missing Modality: A Survey

### v3

@@ 5.1. Data Processing Aspect
5.2. Strategy Design Aspect
(1) Architecture Focused: Some methods focus on designing model training or inference architecture to alleviate the performance degradation caused by missing modalities. Attention-based methods are valued by researchers because they can effectively capture the relationships between various modalities, are scalable in terms of dataset size, and are highly parallelizable. Single attention mechanisms used for modality fusion have the advantage of being plug-and-play. The drawbacks of transformer-based methods in this category are also relatively obvious: training multimodal transformers from scratch requires excessive training time and many GPU resources. Recently popular PEL methods can effectively mitigate this drawback, but the generalization performance of this method is still not as well as full-parameter training or tuning. Although MLLMs can handle an arbitrary number of modalities with flexibility, they are constrained by training complexity and require substantial computational resources.
Distillation-based methods are relatively easy to implement, with student models learning how to reconstruct missing modality representations and inter-modal relationships from teacher models. Since the teacher model typically receives full-modality samples as input, it can simulate any missing modality cases during training, ensuring strong performance. However, most distillation-based methods are limited to the requirement of complete datasets and are inapplicable to datasets with inherent missing modalities. To our knowledge, only one intermediate distillation method (Zhang et al., 2023c) has attempted to input mismatched samples of the same class into the teacher and some self-distillation methods (Wang et al., 2023a; Shi et al., 2024) use representations of available modality branches for distillation purposes.
@@ 5.2. Strategy Design Aspect
Table 2. Another common taxonomy of recovery and non-recovery methods in multimodal learning is based on three stages: early, intermediate, and late stage, as seen in the conventional taxonomy. In this context, recovery and non-recovery methods refer to how missing modalities are handled at each stage.
@@
Modality Generation (Wang et al., 2024a)
Representation Composition (Wang et al., 2023c)
@@
Representation Generation (Ma et al., 2021b)
Distillation-based (Li et al., 2022b)
Coordinated-Representation (Ma et al., 2021c)
@@
Non-Recovery Dedicated (Wang et al., 2021) Discrete Scheduler (Wu et al., 2023b) Representation Composition (Delgado-Escano et al., 2021) Attention-based (Ge et al., 2023) Graph Learning-based (Zhao et al., 2022) Multimodal Large Language Model (Wu et al., 2023a) Coordinated-Representation (Liang et al., 2019) Distillation-based (Shi et al., 2024) Ensemble (Chen et al., 2022) Not Applicable
5.3. Recovery and Non-Recovery Methods & Some Statistics
@@
6. Applications and Datasets
The collection of multimodal datasets is often labor-intensive and costly. In certain specific application directions, issues such as user privacy concerns, sensor malfunctions on data collection devices, and other factors can result in datasets with missing modalities. In severe cases, up to 90% of the samples may have missing modalities, making it challenging for conventional MLFM to achieve good performance. This has given rise to the task of MLMM. Since the factors causing incomplete datasets usually stem from different application directions, we introduce the following datasets based on the applications which is common in MLMM tasks: Sentiment Analysis, Medical Diagnosis, Retrieval/Captioning, Remote Sensing, Robotic Vision, and others. We categorize those datasets according to the applications and data type in Table 3.
Table 3. Common datasets used by deep MLMM methods divided by main applications and modality types. “Vision” includes data from visual sensors such as RGB images, Depth, Infrared, LiDAR, Radar, Event, and Optical Flow. “Bio-Sensors” includes data from sensors like: electrophysiological and respiratory sensors. Other modalities include audio, text, CT scans, MRI, and skeleton.
@@ 6. Applications and Datasets
Sentiment Analysis Vision+Audio eNTERFACE’05 (Martin et al., 2006), RAVDESS (Livingstone and Russo, 2018), CREMA-D (Cao et al., 2014)
@@
Vision+Bio-Sensors Ulm-TSST (Stappen et al., 2021a, b), RECOLA (Ringeval et al., 2013), SEED (Zheng and Lu, 2015)
@@
Medical Diagnosis MRI, CT Scans BRATS-series (Menze et al., 2014; Myronenko, 2019), ADNI (Jack Jr et al., 2008), IXI [https://brain-development.org/ixi-dataset]
Bio-Sensors StressID (Chaptoukaev et al., 2024), PhysioNet (Kemp et al., 2000)
@@
Information Retrieval Vision+Text Amazon Series Datasets (Lakkaraju et al., 2013; McAuley et al., 2015; He and McAuley, 2016; PromptCloud, 2017), MIR-Flickr25K (Duin, [n. d.]), NUS-WIDE (Chua et al., 2009)
@@
Remote Sensing Vision WHU-OPT-SAR (Li et al., 2022c), DFC2020 (Yokoya et al., 2020), MSAW (Shermeyer et al., 2020), Trento (Rasti et al., 2017), Houston (Piacentini, 2023), Berlin (Hong et al., 2021), GRSS (Debes et al., 2014)
Robotic Vision Vision
RegDB (Ye et al., 2021), SYSU-MM01 (Wu et al., 2017), UWA3DII (Rahmani et al., 2016),
Delivery (Zhang et al., 2023a), NuScenes (Caesar et al., 2020), MVSS (Ji et al., 2023)
@@ 6.1. Sentiment Analysis
Typically, the goal of sentiment analysis is to classify the current emotional state by combining information from multiple modalities such as textual, auditory, and RGB information. This field has garnered significant attention due to its promising applications in various specific scenarios, including market research, health monitoring, and advertising. In the early stages of audio-visual sentiment analysis research, researchers discovered that facial detection algorithm in data collection devices sometimes failed to capture faces (e.g., due to occlusion) or the recorded audio is too noisy to use (Cohen et al., 2004; Sebe et al., 2005), resulting in some samples that only contained audio or image. Consequently, researchers began exploring methods to enable sentiment analysis models to continue functioning effectively when facing missing-modality samples. Currently, the datasets for applying sentiment analysis can be roughly divided into two categories. One type of dataset  (Martin et al., 2006; Livingstone and Russo, 2018; Cao et al., 2014; Stappen et al., 2021a, b; Ringeval et al., 2013; Zheng and Lu, 2015; Zadeh et al., 2018; Busso et al., 2008, 2016; Poria et al., 2018; Yu et al., 2020; Wöllmer et al., 2013; Morency et al., 2011) involves using video, audio, text, and biosensors to determine human emotional states in real life or in movies. The other type (Guérin et al., 2013; Nguyen et al., 2018; Matsui et al., 2017) involves using text and images to assess emotional states in contexts such as comic books. Please refer to Table 3 for details.
6.2. Medical Diagnosis
@@
Medical diagnosis requires comprehensive judgments from various modalities such as medical history, physical examination and imaging data, which is exactly what multimodal learning is good at, many researchers developed multimodal intelligent medical diagnosis systems. The current datasets for multimodal medical diagnosis that contain data with missing modality problems are focused on on eight areas: Neuroimaging and Brain Disorders (Menze et al., 2014; Myronenko, 2019; Young et al., 2017; Jack Jr et al., 2008; Lesjak et al., 2018), Cardiovascular (Liu et al., 2016; Moody and Mark, 2001), Cancer Detection (Arya and Saha, 2020; Tomczak et al., 2015), Women Health Analysis (Zhang et al., 2023c; Xu et al., 2021), Ophthalmology (Li et al., 2021; Zhang et al., 2022), Sleep Disorders (Lee et al., 2022; Marcus et al., 2013), Clinical Predictions (Johnson et al., 2016, 2023, 2019), and Biomedical Analysis (Chaptoukaev et al., 2024; Schmidt et al., 2018; Kemp et al., 2000). We have categorized some commonly used and representative datasets from list above by modality type, as shown in Table 3.
6.3. Information Retrieval
@@
Information retrieval is a technology that automatically retrieves relevant content or data based on queries, historical behavior, preferences, and attribute data. It has received widespread attention due to its success on various platforms such as search engines and social media. It improves user satisfaction and information access experience through algorithmic analysis and prediction of content that users may find relevant. Multimodal learning makes modern retrieval systems possible because it can effectively process and analyze multiple types of information, including text, images, audio, etc. Due to user privacy concerns and data sparsity issues, some researchers have begun exploring MLMM approaches in retrieval systems. Many of these studies leverage datasets (Lakkaraju et al., 2013; McAuley et al., 2015; He and McAuley, 2016; PromptCloud, 2017; Duin, [n. d.]; Chua et al., 2009; Xu et al., 2016; Lin et al., 2023) from websites such as Amazon and TikTok for their research. We list those datasets based on modality type in Table 3.
6.4. Remote Sensing
@@
6.5. Robotic Vision
@@
Robotic vision is the field that enables robots to perceive and understand their environment through visual sensors by acquiring and processing image data. This includes tasks such as object and facial recognition, environment modeling, navigation, and behavior understanding, allowing robots to autonomously perform complex operations and interact with humans. Typically, the modality combination formula in robotic vision tasks can be expressed as RGB+X, where X can include, but are not limited to, LiDAR, radar, infrared sensors, depth sensors and auditory sensors. Both RGB and other sensor data can be missing for various reasons. Below, we introduce five common tasks in robotic vision concerning the missing modality problems. (1) Multimodal Segmentation: This task aims to input multimodal data and use segmentation masks to locate objects of interest. In the context of missing modalities, this task is common in datasets for autonomous vehicles, where sensors might fail due to damage or adverse weather conditions. Typical data combinations include RGB+(Depth/Optical Flow/LiDAR/Radar/Infrared/Event data) (Takumi et al., 2017; Liu et al., 2023b; Caesar et al., 2020; Zhang et al., 2023a). Other datasets focus on indoor scene segmentation (Silberman et al., 2012; Ha et al., 2017) (RGB, Depth, Thermal) and material segmentation (Liang et al., 2022) (RGB, Angle of Linear Polarization, Degree of Linear Polarization, Near-Infrared images). (2) Multimodal Detection: Similar to multimodal segmentation, multimodal detection aims to locate objects of interest using bounding boxes. Commonly used multimodal detection combinations in the context of missing modalities include: RGB with Depth (Silberman et al., 2012; Lai et al., 2011) and RGB with Thermal (Hwang et al., 2015; kaggle, 2020). (3) Multimodal Activity Recognition: This task can input data from the aforementioned visual sensors and also use audio cues to recognize human activities. Common datasets (Grauman et al., 2022; Huh et al., 2023; Damen et al., 2022; Gong et al., 2023; Wang et al., 2014; Shahroudy et al., 2016) for this task are combinations of RGB with Audio, Depth, Thermal, Inertial and skeleton data. (4) Multimodal Person Re-Identification: This task aims to use depth and thermal (Ye et al., 2021; Wu et al., 2017) information to robustly identify individuals under varying lighting conditions, such as at night. (5) Multimodal Face Anti-Spoofing: This task utilizes multiple sensors (such as RGB, infrared, and depth cameras) to detect and prevent spoofing in facial recognition systems (John and Kawanishi, 2022; Zhang et al., 2020; Liu et al., 2021b). Its purpose is to enhance the system’s ability to identify fake faces (such as photos, videos, masks, etc.) by comprehensively analyzing different modalities, thereby increasing the security and reliability of facial recognition systems.
@@
Other Applications: There are many other areas where MLMM methods are being explored. For example, conventional audio-visual classification (Vielzeuf et al., 2018); audio-visual question answering (Park et al., 2024); multimodal large language models in visual-dialogue (Bubeck et al., 2023), captioning (Liu et al., 2023a); hand pose estimation using depth images and heat distributions (Tompson et al., 2014; Choi et al., 2017); knowledge graph completion utilizing multimodal data with missing modalities (Liu et al., 2019); multimodal time series prediction, such as stock prediction (Luo et al., 2024) and air quality forecasting (Zheng et al., 2015); multimodal gesture generation on how to create natural animations (Liu et al., 2022); multimodal analysis of single-cell data (Mimitou et al., 2021) in biology.

### v4

@@ 5.1 Data Processing Aspect
5.2 Strategy Design Aspect
(1) Architecture Focused: Some methods focus on designing model training or inference architecture to alleviate the performance degradation caused by missing modalities. Attention-based methods are valued by researchers because they can effectively capture the relationships between various modalities, are scalable in terms of dataset size, and are highly parallelizable. Single attention mechanisms used for modality fusion have the advantage of being plug-and-play. The drawbacks of transformer-based methods in this category are also relatively obvious: training multimodal transformers from scratch requires excessive training time and many GPU resources. Also, those attention-based methods generally require larger datasets due to their high model capacity and reliance on token-level interactions. While this enables stronger representation learning in large-scale settings, it also limits their applicability in data-scarce domains such as medical imaging. This trade-off highlights the need for data-efficient attention mechanisms or pretraining strategies in missing-modality scenarios. Recently popular PEL methods can effectively mitigate this drawback, but the generalization performance of this method is still not as well as full-parameter training or tuning. Although MLLMs can handle an arbitrary number of modalities with flexibility, they are constrained by training complexity and require substantial computational resources.
Distillation-based methods are relatively easy to implement, with student models learning how to reconstruct missing modality representations and inter-modal relationships from teacher models. Since the teacher model typically receives full-modality samples as input, it can simulate any missing modality cases during training, ensuring strong performance. However, most distillation-based methods are limited to the requirement of complete datasets and are inapplicable to datasets with inherent missing modalities. To our knowledge, only one intermediate distillation method (Zhang et al., 2023b) has attempted to input mismatched samples of the same class into the teacher and some self-distillation methods (Wang et al., 2023a; Shi et al., 2024) use representations of available modality branches for distillation purposes.
@@ 5.2 Strategy Design Aspect
5.3 Recovery and Non-Recovery Methods & Some Statistics
Table 2: Another common taxonomy of recovery and non-recovery methods in multimodal learning is based on three stages: early, intermediate, and late stage, as seen in the conventional taxonomy. In this context, recovery and non-recovery methods refer to how missing modalities are handled at each stage.
@@ 5.3 Recovery and Non-Recovery Methods & Some Statistics
Modality Generation (Wang et al., 2024c)
Representation Composition (Wang et al., 2023b)
@@
Representation Generation (Ma et al., 2021c)
Distillation-based (Li et al., 2022c)
Coordinated-Representation (Ma et al., 2021b)
@@
Non-Recovery Dedicated (Wang et al., 2021) Discrete Scheduler (Wu et al., 2023a) Representation Composition (Delgado-Escano et al., 2021) Attention-based (Ge et al., 2023) Graph Learning-based (Zhao et al., 2022) Multimodal Large Language Model (Wu et al., 2023b) Coordinated-Representation (Liang et al., 2019) Distillation-based (Shi et al., 2024) Ensemble (Chen et al., 2022) Not Applicable
@@
Figure 18: The annual statistics are based on our taxonomy of four methods: Modality Imputation, Representation-Focused, Architecture-Focused, and Model Combinations.
5.4 Technical Evolution Discussion
The development of MLMM methods has closely followed broader trends in machine learning and deep learning, with distinct methodological phases reflected in the taxonomy and publication statistics (2014–2025). Based on the papers we collected, we count the papers for each year according to the four major categories in our taxonomy mentioned above, and plotted them using a stacked bar chart (shown in figure 18).
Early research was dominated by data-driven heuristics and ensemble-style solutions, where missing modalities were addressed by retrieving similar samples from the dataset or by training separate unimodal models and combining them through ensemble strategies. This period corresponds to the initial emergence of Modality Imputation and Model Combinations, though the overall scale of research activity was still limited. With the rise of deep learning frameworks (e.g., PyTorch, TensorFlow) and the success of architectures such as ResNet, the community shifted its attention toward representation learning, leading to the rapid growth of Representation-Focused Models. From 2017 onward, this category became the most stable and persistent line of research, as learning aligned, shared, or generative representations enabled models to remain effective even when certain modalities were absent. At the same time, modality-level imputation methods evolved from simple sample-based matching to deep generative models, first through autoencoders and GANs and later through diffusion models, greatly improving the quality and generality of imputed modalities.
A notable transition occurred after 2018, possibly driven by the popularity of attention mechanisms, deep graph neural networks, and knowledge distillation. These techniques allowed models to dynamically adapt to arbitrary combinations of available modalities, leading to the rise of Architecture-Focused Models—a category that grows sharply in the statistics. These methods reframed missing modalities not merely as a data problem but increasingly as an architectural and training-strategy challenge, enabling flexible fusion, conditional routing, and intra/inter-modal knowledge transfer.
Most recently, the emergence of transformer-based foundation models has opened a new direction: applying large pretrained models to missing-modality conditions by using them as feature processors, task schedulers, or backbone architectures capable of consuming variable sets of modalities. This shift substantially expands the design space of multimodal systems and suggests a future where missing-modality robustness is a built-in capability of general-purpose models rather than a specialized add-on.
Finally, although Model Combinations has historically remained a niche line of work, which we suspect is due to the cost of storing and maintaining multiple models, recent interest in Agentic AI systems may renew attention to this class of methods. As agents increasingly integrate multiple specialized tools, dynamic model selection and combination strategies may become more relevant in practical missing-modality scenarios. Together, these developments illustrate a clear technical evolution: from early heuristic data-level fixes to representation learning, to flexible architecture design, and finally toward foundation-model–based general multimodal intelligence. We note that applications and architectures are only weakly linked in this area, largely because modern deep learning architectures (e.g., attention) are highly versatile and transferable across tasks and applications. For instance, generative-adversarial-learning-based recovery methods are used in both breast-cancer prognosis (Arya and Saha, 2021) and visible–infrared person re-identification (Li et al., 2022d), and masked modeling strategy in human action recognition (Woo et al., 2023) closely parallel those used in brain-tumor segmentation (Liu et al., 2023a). As a result, once an architectural idea proves effective for handling missing modalities, it generalizes widely with minimal task-specific adaptation, which naturally weakens the direct coupling between applications and architecture.
@@ 5.4 Technical Evolution Discussion
6 Applications and Datasets
The collection of multimodal datasets is often labor-intensive and costly. In certain specific application directions, issues such as user privacy concerns, sensor malfunctions on data collection devices, and other factors can result in datasets with missing modalities. In severe cases, up to 90% of the samples may have missing modalities, making it challenging for conventional MLFM to achieve good performance. This has given rise to the task of MLMM. Since the factors causing incomplete datasets usually stem from different application directions, we introduce the following datasets based on the applications in MLMM tasks: Affective Computing, Medical Applications, Information Retrieval, Remote Sensing, Pervasive Computing, Robotic Vision and Sensing, and others. We categorize those datasets according to the applications and data type in Table 3. It should be noted that for the medical field, we only show part datasets due to their large number.
Table 3: Common datasets used by deep MLMM methods divided by main applications and modality types. “Vision” includes data from visual sensors such as RGB images, Depth, Infrared, LiDAR, Radar, Event, and Optical Flow. “Bio-Sensors” includes data from sensors like: electrophysiological, respiratory sensors and so on. “Motion Sensors” includes data from sensors like accelerometers, gyroscopes, magnetometers, IMUs, barometric sensors, and gait sensors. “Physiological Signals” including ECG, PPG/BVP, EDA/GSR, RESP, EEG, EMG, EOG, skin temperature, and SpO2. Other modalities include audio, text, CT scans, MRI, and skeleton.
@@ 6 Applications and Datasets
Affective Computing Vision+Audio eNTERFACE’05 (Martin et al., 2006), RAVDESS (Livingstone and Russo, 2018), CREMA-D (Cao et al., 2014)
@@
Vision+Bio-Sensors Ulm-TSST (Stappen et al., 2021a; b), RECOLA (Ringeval et al., 2013), SEED (Zheng and Lu, 2015)
@@
Medical Applications MRI, CT Scans BRATS-series (Menze et al., 2014; Myronenko, 2019), ADNI (Jack Jr et al., 2008), IXI [https://brain-development.org/ixi-dataset]
Bio-Sensors StressID (Chaptoukaev et al., 2024), PhysioNet (Kemp et al., 2000), TCGA [https://portal.gdc.cancer.gov/], BCData (Huang et al., 2020), NuClick (Koohbanani et al., 2020), LYON19 (Swiderska-Chadaj et al., 2019)
@@
Information Retrieval Vision+Text Amazon Series Datasets (Lakkaraju et al., 2013; McAuley et al., 2015; He and McAuley, 2016; PromptCloud, 2017), MIR-Flickr25K (Duin,Robert, ), NUS-WIDE (Chua et al., July 8-10, 2009)
@@
Remote Sensing Vision WHU-OPT-SAR (Li et al., 2022b), DFC2020 (Yokoya et al., 2020), MSAW (Shermeyer et al., 2020), Trento (Rasti et al., 2017), Houston (Piacentini, 2023), Berlin (Hong et al., 2021), GRSS (Debes et al., 2014)
Pervasive Computing Motion sensors Opportunity (Roggen et al., 2010), RealDisp (Banos et al., 2014), DSADS (Barshan and Altun, 2010), PAMAP2 (Reiss, 2012), Shoaib (Shoaib et al., 2014), RealWorld-HAR (Sztyler and Stuckenschmidt, 2016), w-HAR (Burzer et al., 2025), Epilepsy (Villar et al., 2016), EMG Physical Action (Theodoridis, 2011), ERing (Wilhelm et al., 2015), and GENEActiv (Phillips et al., 2013)
Physiological Signals WESAD (Schmidt et al., 2018a), mBrain21 (Donckt, 2024), ETRI Lifelog (Oh et al., 2025)
Electroencephalogram Self RegulationSCP1 (Birbaumer et al., 1999), EEG-MMID (Schalk et al., 2004)
Robotic Vision and Sensing Vision RegDB (Ye et al., 2021), SYSU-MM01 (Wu et al., 2017), UWA3DII (Rahmani et al., 2016), Delivery (Zhang et al., 2023a), NuScenes (Caesar et al., 2020), MVSS (Ji et al., 2023)
@@ 6.1 Affective Computing
Typically, the goal of affective computing is to classify the current emotional state by combining information from multiple modalities such as textual, auditory, and RGB information. This field has garnered significant attention due to its promising applications in various specific scenarios, including market research, health monitoring, and advertising. In the early stages of audio-visual affective computing research, researchers discovered that facial detection algorithm in data collection devices sometimes failed to capture faces (e.g., due to occlusion) or the recorded audio is too noisy to use (Cohen et al., 2004; Sebe et al., 2005), resulting in some samples that only contained audio or image. Consequently, researchers began exploring methods to enable affective computing models to continue functioning effectively when facing missing-modality samples. Currently, the datasets for applying affective computing can be roughly divided into two categories. One type of dataset  (Martin et al., 2006; Livingstone and Russo, 2018; Cao et al., 2014; Stappen et al., 2021a; b; Ringeval et al., 2013; Zheng and Lu, 2015; Zadeh et al., 2018; Busso et al., 2008; 2016; Poria et al., 2018; Yu et al., 2020; Wöllmer et al., 2013; Morency et al., 2011) involves using video, audio, text, and biosensors to determine human emotional states in real life or in movies. The other type (Guérin et al., 2013; Nguyen et al., 2018; Matsui et al., 2017) involves using text and images to assess emotional states in contexts such as comic books. Please refer to Table 3 for details.
6.2 Medical Applications
@@
Medical domain requires comprehensive analysis from various modalities such as medical history, physical examination, imaging data, cell and gene data, which is exactly what multimodal learning is good at, many researchers developed multimodal intelligent systems. The current datasets for multimodal medical domain that contain data with missing modality problems are focused on below areas: Neuroimaging and Brain Disorders (Menze et al., 2014; Myronenko, 2019; Young et al., 2017; Jack Jr et al., 2008; Lesjak et al., 2018), Cardiovascular (Liu et al., 2016; Moody and Mark, 2001), Cancer Detection (Arya and Saha, 2020; Tomczak et al., 2015; Cheerla and Gevaert, 2019; Farooq et al., 2025), Women Health Analysis (Zhang et al., 2023b; Xu et al., 2021), Ophthalmology (Li et al., 2021; Zhang et al., 2022), Sleep Disorders (Lee et al., 2022; Marcus et al., 2013), Clinical Predictions (Johnson et al., 2016; 2023; 2019), Biomedical Analysis (Chaptoukaev et al., 2024; Schmidt et al., 2018b; Kemp et al., 2000), Computational Pathology Cui et al. (2022); Peng et al. (2024); Wang et al. (2025a); Xu et al. (2025), Radiology Chen et al. (2025); Cui et al. (2022); Liang et al. (2025), and Gene Schmauch et al. (2020) and Cell Ghahremani et al. (2022); Liu et al. (2025b); Palma et al. (2024); Tu et al. (2022) Representation learning. We have categorized some commonly used and representative datasets from list above by modality type, as shown in Table 3.
6.3 Information Retrieval
@@
Information retrieval is a technology that automatically retrieves relevant content or data based on queries, historical behavior, preferences, and attribute data. It has received widespread attention due to its success on various platforms such as search engines and social media. It improves user satisfaction and information access experience through algorithmic analysis and prediction of content that users may find relevant. Multimodal learning makes modern retrieval systems possible because it can effectively process and analyze multiple types of information, including text, images, audio, etc. Due to user privacy concerns and data sparsity issues, some researchers have begun exploring MLMM approaches in retrieval systems. Many of these studies leverage datasets (Lakkaraju et al., 2013; McAuley et al., 2015; He and McAuley, 2016; PromptCloud, 2017; Duin,Robert, ; Chua et al., July 8-10, 2009; Xu et al., 2016; Lin et al., 2023) from websites such as Amazon and TikTok for their research. We list those datasets based on modality type in Table 3.
6.4 Remote Sensing
@@
6.5 Pervasive Computing
The pervasive computing community has also developed a wide range of multimodal sensing datasets that capture the complexity of real-world mobile, wearable, and IoT environments. They can be meaningfully organized by task category and reflect the dominant research problems across pervasive and mobile computing. (1) Human Activity Recognition. A broad class of datasets focuses on fine-grained activity recognition, gesture understanding, and daily-living behavior modeling. Representative collections include Opportunity (Roggen et al., 2010), RealDisp (Banos et al., 2014), DSADS (Barshan and Altun, 2010), PAMAP2 (Reiss, 2012), Shoaib (Shoaib et al., 2014), RealWorld-HAR (Sztyler and Stuckenschmidt, 2016), w-HAR (Burzer et al., 2025), ERing (Wilhelm et al., 2015), and EMG Physical Action (Theodoridis, 2011). These datasets capture heterogeneous, body-mounted signals—accelerometry, gyroscope, magnetometer, EMG—collected in naturalistic environments that inherently involve sensor dropout, misalignment, and variable sampling. (2) Physiological Monitoring. Many datasets target long-term physiological monitoring. Examples include WESAD (Schmidt et al., 2018a),Epilepsy (Villar et al., 2016), and combining biosignals such as ECG, EDA, BVP, RESP, or skin temperature, with motion sensors or contextual smartphone streams. Their free-living collection protocol results in multimodal recordings with non-wear intervals, irregular sampling, and missing physiological channels. (3) Cognitive State Inference. Datasets for cognitive-state modeling, mental workload analysis, and motor imagery—such as Self RegulationSCP1 (Birbaumer et al., 1999) and EEG-MMID (Schalk et al., 2004)—contain high-dimensional EEG signals. EEG’s susceptibility to noise, channel corruption, and abrupt signal loss naturally produces structural missingness at both temporal and spatial (channel) levels. Taken together, these task-driven dataset categories illustrate that pervasive computing data are inherently noisy, asynchronous, and incomplete, with many recordings containing long gaps, irregular sampling, and sensor-level failures. As a result, they serve not only as benchmarks for multimodal recognition, context understanding, and mobile health monitoring, but also as natural testbeds for evaluating MLMM methods under real-world structural missingness. Their modality heterogeneity, device-specific characteristics, and episodic sensor dropout further make them valuable for assessing robustness, cross-device generalization, and lightweight recovery strategies necessary for deployment on resource-constrained edge platforms. The classification of above dataset on modalities can be found in Table 3.
6.6 Robotic Vision and Sensing
@@
Robotic vision is the field that enables robots to perceive and understand their environment through visual sensors by acquiring and processing image data. This includes tasks such as object and facial recognition, environment modeling, navigation, and behavior understanding, allowing robots to autonomously perform complex operations and interact with humans. Typically, the modality combination formula in robotic vision tasks can be expressed as RGB+X, where X can include, but are not limited to, LiDAR, radar, infrared sensors, depth sensors and auditory sensors. Both RGB and other sensor data can be missing for various reasons. Below, we introduce five common tasks in robotic vision concerning the missing modality problems. (1) Multimodal Segmentation: This task aims to input multimodal data and use segmentation masks to locate objects of interest. In the context of missing modalities, this task is common in datasets for autonomous vehicles, where sensors might fail due to damage or adverse weather conditions. Typical data combinations include RGB+(Depth/Optical Flow/LiDAR/Radar/Infrared/Event data) (Takumi et al., 2017; Liu et al., 2023b; Caesar et al., 2020; Zhang et al., 2023a). Other datasets focus on indoor scene segmentation (Silberman et al., 2012; Ha et al., 2017) (RGB, Depth, Thermal) and material segmentation (Liang et al., 2022) (RGB, Angle of Linear Polarization, Degree of Linear Polarization, Near-Infrared images). (2) Multimodal Detection: Similar to multimodal segmentation, multimodal detection aims to locate objects of interest using bounding boxes. Commonly used multimodal detection combinations in the context of missing modalities include: RGB with Depth (Silberman et al., 2012; Lai et al., 2011) and RGB with Thermal (Hwang et al., 2015; kaggle, 2020). (3) Multimodal Activity Recognition: This task can input data from the aforementioned visual sensors and also use audio cues to recognize human activities. Common datasets (Grauman et al., 2022; Huh et al., 2023; Damen et al., 2022; Gong et al., 2023; Wang et al., 2014; Shahroudy et al., 2016) for this task are combinations of RGB with Audio, Depth, Thermal, Inertial and skeleton data. (4) Multimodal Person Re-Identification: This task aims to use depth and thermal (Ye et al., 2021; Wu et al., 2017) information to robustly identify individuals under varying lighting conditions, such as at night. (5) Multimodal Face Anti-Spoofing: This task utilizes multiple sensors (such as RGB, infrared, and depth cameras) to detect and prevent spoofing in facial recognition systems (John and Kawanishi, 2022; Zhang et al., 2020; Liu et al., 2021a). Its purpose is to enhance the system’s ability to identify fake faces (such as photos, videos, masks, etc.) by comprehensively analyzing different modalities, thereby increasing the security and reliability of facial recognition systems.
@@
Other Applications: There are many other areas where MLMM methods are being explored. For example, conventional audio-visual classification (Vielzeuf et al., 2018); multimodal large language models in visual-dialogue (Bubeck et al., 2023); audio-visual question answering (Park et al., 2024); captioning (Liu et al., 2023c); hand pose estimation using depth images and heat distributions (Tompson et al., 2014; Choi et al., 2017); knowledge graph completion utilizing multimodal data with missing modalities (Liu et al., 2019); multimodal time series prediction, such as stock prediction (Luo et al., 2024) and air quality forecasting (Zheng et al., 2015); multimodal gesture generation on how to create natural animations (Liu et al., 2022); multimodal analysis of single-cell data (Mimitou et al., 2021) in biology.